# Humanoid Policy $\sim$ Human Policy

### v2

@@ Abstract
Humanoid (adj.): having human form or characteristics. – Merriam-Webster, 2025.
@@
Training manipulation policies for humanoid robots with diverse data enhances their robustness and generalization across tasks and platforms. However, learning solely from robot demonstrations is labor-intensive, requiring expensive tele-operated data collection which is difficult to scale. This paper investigates a more scalable data source, egocentric human demonstrations, to serve as cross-embodiment training data for robot learning. We mitigate the embodiment gap between humanoids and humans from both the data and modeling perspectives. We collect an egocentric task-oriented dataset (PH2D) that is directly aligned with humanoid manipulation demonstrations. We then train a human-humanoid behavior policy, which we term Human Action Transformer (HAT). The state-action space of HAT is unified for both humans and humanoid robots and can be differentiably retargeted to robot actions. Co-trained with smaller-scale robot data, HAT directly models humanoid robots and humans as different embodiments without additional supervision. We show that human data improves both generalization and robustness of HAT with significantly better data collection efficiency.
@@
Figure 1: We propose to use task-oriented egocentric human data to learn humanoid manipulation policies. Without relying on modular perception, we train a Human Action Transformer (HAT) manipulation policy by directly modeling humans as a different humanoid embodiment in an end-to-end manner. To facilitate such cross-embodiment learning, We collect a large-scale dataset, Physical Human-Humanoid Data (PH2D), with automatic hand-finger 3D poses from consumer-grade VR devices.
I Introduction
@@
Learning from real robot demonstrations has led to great progress in robotic manipulation recently [34, 40, 6, 15]. One key advancement to enable such progress was hardware / software co-designs to scale up data collection using teleoperation [65, 20, 11, 58, 9, 25] and directly controlling the robot end effector [14, 5, 65, 20, 23, 11]. Instead of gathering data on a single robot, collective efforts have been made to merge diverse robot data and train foundational policies across embodiments [14, 41, 40, 34, 6, 15], which have shown to improve cross-embodiment and cross-task generalizability. However, collecting structured real-robot data is expensive and time-consuming. We are still far away from building a robust and generalizable model as what has been achieved in Computer Vision [46] and NLP [42].
If we examine humanoid robot teleoperation more closely, it involves robots mimicking human actions using geometric transforms or retargeting to control robot joints and end-effectors. From this perspective, we propose to model robots in a human-centric representation, and the robot action is just a transformation away from the human action. If we can accurately capture the end-effector and head poses of humans, egocentric human demonstrations will be a more scalable source of training data, as we can collect them efficiently, in any place, and without a robot.
This paper performs cross-human and humanoid embodiment training for robotic manipulation. Our key insight is to model bimanual humanoid behaviors by directly imitating human behaviors without using learning surrogates such as affordances [37, 3]. To realize this, we first collect an egocentric task-oriented dataset of Physical Humanoid-Human Data, dubbed PH2D. We adapt consumer-grade VR devices to collect egocentric videos with automatic but accurate hand pose and end effector (i.e., hand) annotations. Compared to existing human daily behavior datasets [22, 12], PH2D is task-oriented so that it can be directly used for co-training. The same VR hardwares are then used to perform teleoperation to collect smaller-scale humanoid data for better alignment. We then train a Human-humanoid Action Transformer (HAT), which predicts future hand-finger trajectories in a unified human-centric state-action representation space. To obtain robot actions, we simply apply inverse kinematics and hand retargeting to convert human actions to robot actions. Such a conversion is a differentiable process that allows end-to-end training on different embodiments.
We conduct real-robot evaluations on different manipulation tasks with extensive ablation studies to investigate how to best align human and humanoid demonstrations. In particular, we found that co-training with diverse human data improves robustness against spatial variance and background perturbation. In addition, it enables generalization to objects completely unseen in robot data. We believe that these findings highlight the potential of using human data for large-scale cross-embodiment learning.
Dataset Human Robot
# Frames # Demos # Frames # Demos
DexCap [51] $\sim$378k 787 NA NA
EgoMimic [28] $\sim$432k 2,150 1.29M 1,000
PH2D (Ours) $\sim$3.02M 26,824 $\sim$668k 1,552
@@
TABLE I: Comparisons of task-oriented egocentric human datasets. Besides having the most demonstrations, PH2D is collected on various manipulation tasks, a diverse set of objects and receptacles, accompanied by accurate 3D hand-finger poses and language annotations. The number of frames is estimated assuming 30 Hz. : estimated based on time reported for data collection; whereas DexCap [51] and PH2D report frames after initial processing.
@@
  • A dataset, PH2D, which is a large egocentric, task-oriented human-humanoid dataset with accurate hand and wrist poses for modeling human behavior (see Tab. I).
@@
II Related Work
Imitation Learning for Robot Manipulation. Recently, learning robot policy with data gathered directly from the multiple and target robot embodiment has shown impressive robustness and dexterity [66, 40, 52, 34, 10, 45, 9, 35]. The scale of data for imitation learning has grown substantially with recent advancements in data collection [1, 9, 11, 58], where human operators can efficiently collect large amounts of high-quality, task-oriented data. Despite these advances, achieving open-world generalization still remains a significant challenge due to lack of internet-scale training data.
@@ II Related Work
II-1 Aligning Observations
While teleoperating the actual robot platform allows learning policy with great dexterity, there is still a long way to go to achieve higher levels of generalization across diverse tasks, environments, and platforms. Unlike fields such as computer vision [46] and natural language processing [42] benefiting from internet-scale data, robot data collection in the real world is far more constrained. Various approaches have attempted to use internet-scale human videos to train robot policies [7, 30, 31, 39, 48, 59]. Due to various discrepancies (e.g., supervision and viewpoints) between egocentric robot views and internet videos, most existing work [37, 3] use modular approaches with intermediate representations as surrogates for training. The most representative ones are affordances [37, 3] for object interaction, object keypoints predictions [4, 53, 32, 13, 54], or other types of object representations [44, 38, 36].
II-2 Aligning Actions
Beyond observation alignment, transferring human demonstrations to robotic platforms introduces additional challenges due to differences in embodiment, actuation, and control dynamics. Specific alignment of human and robot actions is required to overcome these disparities. Approaches have employed masking in egocentric views [28], aligning motion trajectories or flow [33, 47], object-centric actions [69, 26], or hand tracking with specialized hardware [51].
However, many existing papers focus on imitating humans in a single task, overlooking the potential of human data to be directly applied for larger scale cross-embodiment learning.
@@ II-2 Aligning Actions
Concurrent Work. Some concurrent work [51, 28, 50] also attempts to use egocentric human demonstrations for end-to-end cross-embodiment policy learning. DexCap [51] uses gloves to track 3D hand poses with a chest-mounted RGBD camera to capture egocentric human videos. However, DexCap relies on 3D inputs, whereas some recent works [6, 34] have shown the scalability of 2D visual inputs. Most related to our work, EgoMimic [28] also proposes to collect data using wearable device [17] with 2D visual inputs. However, EgoMimic requires strict visual sensor alignment and heuristic designs such as visual masking. Such reliance on visual models during training and testing leads to composite failure similar to the modular approach. In addition, PH2D is also greater in dataset scale and object diversity. We also show our policy can be deployed on real robots without strict requirements of visual sensors and heuristics, which paves the way for scalable data collection.
III Method
To collect more data to train generalizable robot policies, recent research has explored cross-embodiment learning, enabling policies to generalize across diverse physical forms [6, 34, 15, 40, 29, 41]. This paper proposes egocentric human manipulation demonstrations as a scalable source of cross-embodiment training data. Sec. III-A describes our approach to adapt consumer-grade VR devices to scale up human data collection conveniently for a dataset of task-oriented egocentric human demonstrations. Sec. III-B describes various techniques to handle domain gaps to align human data and robot data for learning humanoid manipulation policy.
III-A PH2D: Task-oriented Physical Humanoid-Human Data
Though there has been existing work that collects egocentric human videos [28, 12, 22, 51], they either (1) provide demonstrations mostly for non-task-oriented skills and do not provide world-frame 3D head and hand poses estimations for imitation learning supervision [22, 12] or (2) require specialized hardware or robot setups [51, 28].
@@ III-A PH2D: Task-oriented Physical Humanoid-Human Data
Figure 2: Adapting Consumer-grade Devices for Data Collection. To avoid relying on specialized hardware for data collection and make our method more accessible, we design our data collection process using consumer-grade VR devices.
@@
With development in pose estimation [68] and system engineering, modern mobile devices are capable of providing accurate on-device world frame 3D head pose tracking and 3D hand keypoint tracking [9], which has proved to be stable enough to teleoperate robot in real-time [9, 23]. We design software and hardware to support convenient data collection across different devices:
  • Apple Vision Pro + Built-in Camera. We developed a Vision OS App that accesses the bottom left camera for visual observation and uses the Apple ARKit to access 3D head and hand poses.
  • Meta Quest 3 / Apple Vision Pro + ZED Camera. We developed a web-based application based on OpenTelevision [9] to gather 3D head and hand poses. We also designed a 3D-printed holder to mount ZED Mini Stereo cameras on these devices. This configuration is both low-cost ($<$700\$) and introduces more diversity with stereo cameras.
Different cameras are intended to provide more visual diversity for data collection.
@@
Action Domain Gap. Human actions and tele-operated robot actions exhibit two distinct characteristics: (1) human manipulation usually involves involuntary whole-body movement, and (2) humans are more dexterous than robots and have significantly faster task completion time than robots. We mitigate the first gap by requesting the human data collectors to sit in an upright position. In addition, we place objects close to humans so that they are within the workspace of human arms without whole-body movement, which resembles the workspace of commercial humanoid robots without whole-body movement.
III-B HAT: Human Action Transformer
@@
More concretely, let $\mathcal{D}_{robot}=\{(\mathbf{S}_{i},\mathbf{A}_{i})\}_{i=1}^{N}$ be the set of data collected from real bimanual humanoid robots using teleoperation [9], where $\mathbf{S}_{i}$ is the states including proprioceptive and visual observations of $i$-th demonstration and $\mathbf{A}_{i}$ be the actions. The collected PH2D dataset, $\mathcal{D}_{human}=\{(\mathbf{\tilde{S}}_{i},\mathbf{\tilde{A}}_{i})\}_{i=1}^{M}$ is used to augment the training process. Note that it is reasonable to assume $M\gg N$ because collecting egocentric human videos can be done more efficiently than both tele-operation.
@@
The goal is to design a policy $\pi:\mathbf{S}\to\mathbf{A}$ that predicts future robot actions $\mathbf{a}_{t}$ given current robot observation $\mathbf{s}_{t}$ at time $t$, where the future actions $\mathbf{a}_{t+1}$ is usually a chunk of actions for multi-step execution (with slight abuse of notation). This paper uses HAT as the policy, which is based on Action Chunk Transformer [65]. We modified the original implementation to replace ResNet-18 [24] with frozen DinoV2 [43] backbones and added visual adaptor to intermediate layers. The overview of the model is illustrated in Fig. 3. We discuss key design choices of HAT with experimental ablations.
Unified State-Action Space. Both bimanual robots and humans have two end effectors. In our case, our robot is also equipped with an actuated 2DoF neck that can rotate, which resembles the autonomous whole-body movement when humans perform manipulation. Therefore, we design a unified state-action space (i.e., $(\mathbf{S},\mathbf{A})\equiv(\mathbf{\tilde{S}},\mathbf{\tilde{A}})$) for both bimanual robots and humans. More concretely, the action space for prediction of HAT is a 54-dimensional vector. Rotations of the head, left wrist, and right wrist are represented as 6D rotations [67]; the translation of left and right wrists are represented as x/y/z vectors. In this work, since we deploy our policy on a robot with 5-fingered dexterous hands (shown in Fig. 3), there exists a bijective mapping between 10 finger tips of robot dexterous hands and common human hands, which are represented as 3D x/y/z keypoints for regressing. Note that injective mapping is also possible (e.g., mapping distance between the thumb finger and other fingers to parallel gripper distance, but we leave it for future study).
@@
Visual Domain Gap. Two types of domain gaps exist for co-training on human/humanoid data: camera sensors and end effector appearance. Since our human data collection process includes cameras different from robot deployment, this leads to camera domain gaps such as tones. In addition, the appearances of human and humanoid end effectors are different. However, with sufficiently large and diverse data, we find it not a strict necessity to apply heuristic policies such as adding visual artifacts [28] or generative methods [61] to train human-robot policies - basic image augmentations such as color jittering and Gaussian blurring are effective regularization of the visual inputs in this case.
Action Domain Gap. To mitigate the drastic speed difference between humans and humanoids, we interpolate translation and rotations of human data during training (effectively ‘slowing down’ actions). The slow-down factors $\alpha_{\text{slow}}$ are obtained by normalizing the average task completion time of humans and humanoids, which is empirically distributed around $4$. For consistency, we use $\alpha_{\text{slow}}=4$ in all tasks.
Due to the difference in state space, we randomly dropout proprioception readings during training to avoid undesired reliance on low-dimensional state inputs.
(a) The robot performs the cup passing task across four different backgrounds. The left side shows the four background variations, while the right side illustrates the two passing directions: (#1 - Right hand passes the cup to the left hand, #2 - Left hand passes the cup to the right hand).
(b) The robot performs the horizontal grasping task with four different items: bottle, box_1, box_2, and can, as shown on the left. The right side illustrates the process: (#1-#3 - The robot grasps the bottle, #4-#5 - The robot places it into the plastic bin).
(c) The robot performs the vertical grasping task. As shown on the left, the Dynamixel box is placed in nine different positions for grasping. The right side illustrates the process: (#1-#3 - The robot grasps the box, #4-#5 - The robot places the box into the plastic bin).
(d) The robot performs the pouring task. The left side shows different settings achieved by varying the robot’s rotation and the table’s position. The right side illustrates the pouring process: (#1 - Right hand grasps the bottle, #2 - Left hand grasps the cup, #3 - Pouring the drink, #4 - Left hand places the cup down, #5 - Right hand places the bottle down).
Figure 4: Illustrations of tasks used in quantitative evaluations. From top to bottom: cup passing, horizontal grasping, vertical grasping, and pouring.
@@
Training. The final policy is denoted as $\pi:f_{\theta}(\cdot)\to\mathbf{A}$ for both human and robot policy. The trainable parameters of both the proprioceptive encoder $\theta$ and the transformer trunks are optimized jointly during training. The final loss is given by,
@@
where $\pi(s_{i})_{\text{EEF}}$ and $a_{i,\text{EEF}}$ are predicted 6dof end-effector poses to emphasize the importance of end effector positions over learning unnecessarily precise finger tip keypoints, $\lambda=2$ is an (insensitive) hyperparameter used to balance loss.
IV Experiments

### v3

@@ Abstract
Training manipulation policies for humanoid robots with diverse data enhances their robustness and generalization across tasks and platforms. However, learning solely from robot demonstrations is labor-intensive, requiring expensive tele-operated data collection, which is difficult to scale. This paper investigates a more scalable data source, egocentric human demonstrations, to serve as cross-embodiment training data for robot learning. We mitigate the embodiment gap between humanoids and humans from both the data and modeling perspectives. We collect an egocentric task-oriented dataset (PH2D) that is directly aligned with humanoid manipulation demonstrations. We then train a human-humanoid behavior policy, which we term Human Action Transformer (HAT). The state-action space of HAT is unified for both humans and humanoid robots and can be differentiably retargeted to robot actions. Co-trained with smaller-scale robot data, HAT directly models humanoid robots and humans as different embodiments without additional supervision. We show that human data improve both generalization and robustness of HAT with significantly better data collection efficiency.
@@
Figure 1: This paper advocates high-quality human data as a data source for cross-embodiment learning - task-oriented egocentric human data. We collect a large-scale dataset, Physical Human-Humanoid Data (PH2D), with hand-finger 3D poses from consumer-grade VR devices on well-defined manipulation tasks directly aligned with robots. Without relying on modular perception, we train a Human Action Transformer (HAT) manipulation policy by directly modeling humans as a different humanoid embodiment in an end-to-end manner.
Keywords: Robot Manipulation, Cross-Embodiment, Humanoid
1 Introduction
@@
Learning from real robot demonstrations has led to great progress in robotic manipulation recently [1, 2, 3, 4]. One key advancement to enable such progress was hardware / software co-designs to scale up data collection using teleoperation [5, 6, 7, 8, 9, 10] and directly controlling the robot end effector [11, 12, 5, 6, 13, 7]. Instead of gathering data on a single robot, collective efforts have been made to merge diverse robot data and train foundational policies across embodiments [11, 14, 2, 1, 3, 4], which have shown to improve cross-embodiment and cross-task generalizability.
@@
Table 1: Comparisons of task-oriented egocentric human datasets. Besides having the most demonstrations, PH2D is collected on various manipulation tasks, diverse objects and scenes, with accurate 3D hand-finger poses and language annotations. : estimated based on reported data collection time with 30 Hz; whereas DexCap [15] and PH2D report processed frames for training.
However, collecting structured real-robot data is expensive and time-consuming. We are still far away from building a robust and generalizable model as what has been achieved in Computer Vision [17] and NLP [18]. If we examine humanoid robot teleoperation more closely, it involves robots mimicking human actions using geometric transforms or retargeting to control robot joints and end-effectors. From this perspective, we propose to model robots in a human-centric representation, and the robot action is just a transformation away from the human action. If we can accurately capture the end-effector and head poses of humans, egocentric human demonstrations will be a more scalable source of training data, as we can collect them efficiently, in any place, and without a robot.
In this paper, we perform cross-human and humanoid embodiment training for robotic manipulation. Our key insight is to model bimanual humanoid behaviors by directly imitating human behaviors without using learning surrogates such as affordances [19, 20]. To realize this, we first collect an egocentric task-oriented dataset of Physical Humanoid-Human Data, dubbed PH2D. We adapt consumer-grade VR devices to collect egocentric videos with automatic but accurate hand pose and end effector (i.e., hand) annotations. Compared to existing human daily behavior datasets [21, 22], PH2D is task-oriented so that it can be directly used for co-training. The same VR hardwares are then used to perform teleoperation to collect smaller-scale humanoid data for better alignment. We then train a Human-humanoid Action Transformer (HAT), which predicts future hand-finger trajectories in a unified human-centric state-action representation space. To obtain robot actions, we simply apply inverse kinematics and hand retargeting to differentiably convert human actions to robot actions for deployment.
We conduct real-robot evaluations on different manipulation tasks with extensive ablation studies to investigate how to best align human and humanoid demonstrations. In particular, we found that co-training with diverse human data improves robustness against spatial variance and background perturbation, generalizing in settings unseen in robot data but seen in human data. We believe that these findings highlight the potential of using human data for large-scale cross-embodiment learning.
@@
  • A dataset, PH2D, which is a large egocentric, task-oriented human-humanoid dataset with accurate hand and wrist poses for modeling human behavior (see Tab. 1).
@@
2 Related Work
Imitation Learning for Robot Manipulation. Recently, learning robot policy with data gathered directly from the multiple and target robot embodiment has shown impressive robustness and dexterity [23, 2, 24, 1, 25, 26, 9, 27, 28]. The scale of data for imitation learning has grown substantially with recent advancements in data collection [29, 9, 7, 8], where human operators can efficiently collect large amounts of high-quality, task-oriented data. Despite these advances, achieving open-world generalization still remains a significant challenge due to lack of internet-scale training data.
@@ 2 Related Work
Learn from Human - Aligning Observations. While teleoperating the actual robot platform allows learning policy with great dexterity, there is still a long way to go to achieve higher levels of generalization across diverse tasks, environments, and platforms. Unlike fields such as computer vision [17] and natural language processing [18] benefiting from internet-scale data, robot data collection in the real world is far more constrained. Various approaches have attempted to use internet-scale human videos to train robot policies [30, 31, 32, 33, 34, 35]. Due to various discrepancies (e.g., supervision and viewpoints) between egocentric robot views and internet videos, most existing work [19, 20] use modular approaches with intermediate representations as surrogates for training. The most representative ones are affordances [19, 20] for object interaction, object keypoints predictions [36, 37, 38, 39, 40], or other types of object representations [41, 42, 43].
Learn from Human - Aligning Actions. Beyond observation alignment, transferring human demonstrations to robotic platforms introduces additional challenges due to differences in embodiment, actuation, and control dynamics. Specific alignment of human and robot actions is required to overcome these disparities. Approaches have employed masking in egocentric views [16], aligning motion trajectories or flow [44, 45], object-centric actions [46, 47], or hand tracking with specialized hardware [15]. Most closely related to our work, HumanPlus [48] designs a remapping method from 3D human pose estimation to tele-operate humanoid robots. Compared to HumanPlus, the insight of our method is to waive the requirement for robot hardware in collecting human data and collect diverse human data directly for co-training. In contrast to HumanPlus, we intentionally avoid performing retargeting on human demonstrations and designed the policy to directly use human hand poses as states/actions. On the other hand, the ‘human shadowing’ retargeting in HumanPlus is a teleoperation method that still requires robots, leading to lower collection efficiency than ours.
@@
Concurrent Work. Some concurrent work [15, 16, 63] also attempts to use egocentric human demonstrations for end-to-end cross-embodiment policy learning. DexCap [15] uses gloves to track 3D hand poses with a chest-mounted RGBD camera to capture egocentric human videos. However, DexCap relies on 3D inputs, whereas some recent works [3, 1] have shown the scalability of 2D visual inputs. Most related to our work, EgoMimic [16] also proposes to collect data using wearable device [64] with 2D visual inputs. However, EgoMimic requires strict visual sensor alignments; whereas we show that scaling up diverse observations with different cameras makes the policy more robust. In addition, PH2D is also greater in dataset scale and object diversity. We also show our policy can be deployed on real robots without strict requirements of visual sensors and heuristics, which paves the way for scalable data collection.
3 Method
To collect more data to train generalizable robot policies, recent research has explored cross-embodiment learning, enabling policies to generalize across diverse physical forms [3, 1, 4, 2, 65, 14]. This paper proposes egocentric human manipulation demonstrations as a scalable source of cross-embodiment training data. Sec. 3.1 describes our approach to adapt consumer-grade VR devices to scale up human data collection conveniently for a dataset of task-oriented egocentric human demonstrations. Sec. 3.2 describes various techniques to handle domain gaps to align human data and robot data for learning humanoid manipulation policy.
3.1 PH2D: Task-oriented Physical Humanoid-Human Data
Though there has been existing work that collects egocentric human videos [16, 22, 21, 15], they either (1) provide demonstrations mostly for non-task-oriented skills (e.g., dancing) and do not provide world-frame 3D head and hand poses estimations for imitation learning supervision [21, 22] or (2) require specialized hardware or robot setups [15, 16].
Figure 2: Consumer-grade Devices for Data Collection. To avoid relying on specialized hardware for data collection to make our method scalable, we design our data collection process using consumer-grade VR devices.
@@ 3.1 PH2D: Task-oriented Physical Humanoid-Human Data
With development in pose estimation [66] and system engineering, modern mobile devices are capable of providing accurate on-device world frame 3D head pose tracking and 3D hand keypoint tracking [9], which has proved to be stable enough to teleoperate robot in real-time [9, 13]. We design software and hardware to support convenient data collection across different devices. Different cameras provide better visual diversity.
  • Apple Vision Pro + Built-in Camera. We developed a Vision OS App that uses the built-in camera for visual observation and uses the Apple ARKit for 3D head and hand poses.
  • Meta Quest 3 / Apple Vision Pro + ZED Camera. We developed a web-based application based on OpenTelevision [9] to gather 3D head and hand poses. We also designed a 3D-printed holder to mount ZED Mini Stereo cameras on these devices. This configuration is both low-cost (<<700\$) and introduces more diversity with stereo cameras.
@@
Action Domain Gap. Human actions and tele-operated robot actions exhibit two distinct characteristics: (1) human manipulation usually involves involuntary whole-body movement, and (2) humans are more dexterous than robots and have significantly faster task completion time than robots. We mitigate the first gap by requesting the human data collectors to sit in an upright position. For the second speed gap, we interpolate translation and rotations of human data during training (effectively ‘slowing down’ actions). The slow-down factors $\alpha_{\text{slow}}$ are obtained by normalizing the average task completion time of humans and humanoids, which is empirically distributed around $4$. For consistency, we use $\alpha_{\text{slow}}=4$ in all tasks.
3.2 HAT: Human Action Transformer
@@
More concretely, let $\mathcal{D}_{robot}=\{(\mathbf{S}_{i},\mathbf{A}_{i})\}_{i=1}^{N}$ be the set of data collected from real bimanual humanoid robots using teleoperation [9], where $\mathbf{S}_{i}$ is the states including proprioceptive and visual observations of $i$-th demonstration and $\mathbf{A}_{i}$ be the actions. The collected PH2D dataset, $\mathcal{D}_{human}=\{(\mathbf{\tilde{S}}_{i},\mathbf{\tilde{A}}_{i})\}_{i=1}^{M}$ is used to augment the training process. Note that it is reasonable to assume $M\gg N$ due to the significantly better human data collection efficiency.
@@
The goal is to design a policy $\pi:\mathbf{S}\to\mathbf{A}$ that predicts future robot actions $\mathbf{a}_{t}$ given current robot observation $\mathbf{s}_{t}$ at time $t$, where the future actions $\mathbf{a}_{t+1}$ is usually a chunk of actions for multi-step execution (with slight abuse of notation). We model $\pi$ using HAT, which is a transformer-based architecture predicting action chunks [5]. The overview of the model is illustrated in Fig. 3. We discuss key design choices of HAT with experimental ablations.
[…]
@@
Visual Domain Gap. Two types of domain gaps exist for co-training on human/humanoid data: camera sensors and end effector appearance. Since our human data collection process includes cameras different from robot deployment, this leads to camera domain gaps such as tones. Also, the appearances of human and humanoid end effectors are different. However, with sufficiently large and diverse data, we find it not a strict necessity to apply heuristic processing such as visual artifacts [16] or generative methods [69] to train human-robot policies - basic image augmentations such as color jittering and Gaussian blurring are effective regularization.
@@
Training. The final policy is denoted as $\pi:f_{\theta}(\cdot)\to\mathbf{A}$ for both human and robot policy, where $f_{\theta}$ is a transformer-based neural network parametrized by $\theta$. The final loss is given by,
@@
where EEF are the indices of the translation vectors of the left and right wrists, and $\lambda=2$ is an (insensitive) hyperparameter used to balance loss to emphasize the importance of end effector positions over learning unnecessarily precise finger tip keypoints.
4 Experiments